# Insights into the Lottery Ticket Hypothesis and Iterative Magnitude Pruning

### v2

@@ IV Results and Findings
IV-B Result 2: There exists a barrier between IMP solutions at successive levels in the loss landscape.
@@
One interpretation of the previous result would clearly be that IMP solutions at different levels are distinct minima. In order to test this hypothesis more definitively, we examine whether these are separated by distinct barriers in the loss landscape. If this was not the case, SGD should converge to a sparser solution in the first place and would not stop at a less sparse solution. Evidence for our hypothesis is presented in Fig. 9. It gives the training loss along a straight line connecting $W^{(min\_(L-1))}_{(L-1)}$ and $W^{(min\_(L))}_{(L)}$. The barriers between successive minima are clearly visible. Thus, at least in the loss landscape, the IMP solutions at successive levels are not linearly connected, unlike the corresponding claim by [17] in the test error landscape. The implication is that these loss barriers are large enough to prevent SGD from crossing over and small enough to have almost similar generalization performance all over the region between these solutions.
@@
Figure 9: Training Loss along a straight line connecting $W^{(min\_(L-1))}_{(L-1)}$ and $W^{(min\_(L))}_{(L)}$. The x-axis represents the interpolation co-efficient $\alpha$. Each plot depicts the training loss at $501$ points between $W^{(min\_(L-1))}_{(L-1)}$ and $W^{(min\_(L))}_{(L)}$.
@@
IV-C Result 3: IMP solutions obtained using rewinding lie within the same loss sublevel set.
@@
Any random initialization of a dense neural network makes SGD converge to a good minimum. This can be ascribed to the well-understood fact [26] that the density of good minima in such scenarios is very high. However, in sparse subspaces, the density of such minima is much smaller. If we consider starting from a random point in such subspaces, the solution would not be as good as we get with rewinding, as is observed in our experiments.
@@
We also hypothesize that all the IMP solutions obtained using rewinding lie within the same loss sublevel set as that of the original dense network, whereas random initialization of the pruned network takes SGD out of that sublevel set. Fig. 10 presents the evidence for our claim. The figure clearly shows that all the IMP solutions ($W^{(min\_(0))}_{(0)}$ to $W^{(min\_(10))}_{(10)}$) lie within the same connected loss sublevel set (dark green region). However, $W^{(RIPN)}$ lies outside the sublevel set. Similarly, random pruning of a large number of weights also takes SGD outside the sublevel set. It is interesting to see that $W^{(one\_shot)}$ also lies in the same sublevel set as the IMP solutions; however, $W^{(RPN\_2)}$ lies outside the sublevel set. In other words, a randomly pruned network obtained by pruning a large number of parameters does not yield an equally good solution as that of the network obtained with magnitude based pruning, as it is not seen to lie in the same sublevel set.
@@
Figure 11: Euclidean distance between different points of interest. Left: Euclidean distance between IMP solutions at different levels. Right: Euclidean distance of $W^{(RIPN)}$, $W^{(RPN\_1)}$, $W^{(RPN\_2)}$, $W^{(FT)}$ and $W^{(one\_shot)}$ with $W^{(min\_(0))}_{(0)}$, $W^{(min\_(9))}_{(9)}$ and $W^{(min\_(10))}_{(10)}$
Figure 12: Cosine similarity between different points of interest. Left: Cosine similarity between IMP solutions at different levels. Right: Cosine similarity of $W^{(RIPN)}$, $W^{(RPN\_1)}$, $W^{(RPN\_2)}$, $W^{(FT)}$ and $W^{(one\_shot)}$ with $W^{(min\_(0))}_{(0)}$, $W^{(min\_(9))}_{(9)}$ and $W^{(min\_(10))}_{(10)}$
It is apparent from these figures that the euclidean distance between IMP solutions is smaller than the euclidean distance of $W^{(RIPN)}$ with $W^{(min\_(9))}_{(9)}$ and $W^{(min\_(10))}_{(10)}$. Similarly, the cosine similarity between IMP solutions is greater than the cosine similarity of $W^{(RIPN)}$ with $W^{(min\_(9))}_{(9)}$ and $W^{(min\_(10))}_{(10)}$. Also the euclidean distance between $W^{(min\_(0))}_{(0)}$ and $W^{(one\_shot)}$ is smaller than the euclidean distance between $W^{(min\_(0))}_{(0)}$ and $W^{(RPN\_2)}$, while the cosine similarity between $W^{(min\_(0))}_{(0)}$ and $W^{(one\_shot)}$ is greater than the cosine similarity of $W^{(min\_(0))}_{(0)}$ with $W^{(RPN\_2)}$. This further buttresses the claim.
IV-D Result 4: Why does the initialization proposed by the lottery ticket hypothesis work well?
The above results also provide a strong clue to the positive role of the initialization proposed by the lottery ticket hypothesis in finding good minima in the sparser weight space. As seen so far, pruning out the smaller weights leads to a new minimum or a saddle point with a slightly larger training loss value and a smaller volume. Retraining from this point via learning rate rewinding or from a point in close proximity (via weight rewinding) leads SGD to converge to a new minimum, which has a larger volume (compared to the earlier minimum) and a lower training loss. This minimum was not discoverable earlier since it spanned some dimensions where the loss function was steeply increasing and therefore, had an overall smaller volume, and pruning out these dimensions exposed this minimum to the SGD. Hence, the starting minimum acts as a baseline, avoidance of which leads to a better minimum. Choosing any other initialization point places the SGD at a point outside the loss sublevel set and there is no guarantee of finding a better minimum. Fig. 13 presents the training loss along a straight line connecting $W^{(min\_(9))}_{(9)}$ (baseline for $W^{(RIPN)}$) and $W^{(RIPN)}$.
Figure 13: Training loss along a straight line between $W^{(min\_(9))}_{(9)}$ and $W^{(RIPN)}$.
Further, a comparison of top-100 positive eigen values of the Hessian at $W^{(RIPN)}$ and $W^{(min\_(10))}_{(10)}$ (Fig. 14) shows that $W^{(RIPN)}$ has larger eigen values than $W^{(min\_(10))}_{(10)}$, which indicates a smaller volume for the basin around $W^{(RIPN)}$ (Table V).
Figure 14: Comparison of top-100 positive eigen values of the Hessian at $W^{(RIPN)}$ and $W^{(min\_(10))}_{(10)}$.
TABLE V: Comparison of $\lambda_{max}$ and $V^{{}^{\prime}}(100)$ at $W^{(RIPN)}$ and $W^{(min-(10))}_{(10)}$.
Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
$W^{(RIPN)}$ 6271.873 497.099
$W^{(min\_(10))}_{(10)}$ 1575.910 408.810
IV-E Result 5: Why do we need the Iterative Process, and why does one-shot pruning not work as well?
A comparison of top-100 positive eigen values of the Hessian between $W^{(one\_shot)}$ and $W^{(min\_(10))}_{(10)}$ (Fig. 15) shows that $W^{(one\_shot)}$ has larger eigen values than $W^{(min\_(10))}_{(10)}$, which implies a smaller volume for the basin around $W^{(one\_shot)}$ (Table VI). This brings out and confirms the importance of the iterative process in pruning.
Figure 15: Comparison of top-100 positive eigen values of the Hessian at $W^{(one\_shot)}$ and $W^{(min\_(10))}_{(10)}$.
TABLE VI: Comparison of $\lambda_{max}$ and $V^{{}^{\prime}}(100)$ at $W^{(one\_shot)}$ and $W^{(min\_(10))}_{(10)}$
Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
$W^{(one\_shot)}$ 2944.881 450.409
$W^{(min\_(10))}_{(10)}$ 1575.910 408.810
@@ IV-F Result 6: What happens when you prune the smaller weights?
$Loss({W_{pruned}})=Loss({W})+{(W_{pruned}-W)^{T}}\frac{\partial Loss}{\partial{W}}_{{W}={W}}+{(W_{pruned}-W)^{T}}\frac{\partial^{2}Loss}{\partial{W}^{2}}{(W_{pruned}-W)}$ (5)
@@
Pruning out the smallest components in the weight vector ${W}$ results in only a small increase in the loss value. Pruning out the larger weights tends to bring about a bigger increase in the loss value since it makes the vector ${(W_{pruned}-W)}$ have a larger magnitude (norm). This can invalidate the second-order approximation given in Eq. $5$, since, in general, the loss function will not be a smooth one, and may have several points of inflection in the neighborhood.
@@
Randomly pruning a large number of weights also has a similar effect in that it may lead SGD to a position outside the sublevel set region, so that even re-training with weight rewinding cannot bring the network back to the sublevel set. Fig. 17 presents the training loss along a straight line connecting $W^{(min\_(0))}_{(0)}$ and $W^{(one\_shot)}$, and $W^{(min\_(0))}_{(0)}$ (baseline for $W^{(RPN\_2)}$) and $W^{(RPN\_2)}$.
Figure 17: Left: Training Loss along a straight connecting $W^{(min\_(0))}_{(0)}$ and $W^{(one\_shot)}$. Right: Training Loss along a straight connecting $W^{(min\_(0))}_{(0)}$ and $W^{(RPN\_2)}$.
The figure shows a small barrier between $W^{(min\_(0))}_{(0)}$ and $W^{(one\_shot)}$, and a huge barrier between $W^{(min\_(0))}_{(0)}$ and $W^{(RPN\_2)}$, which again demonstrates that $W^{(min\_(0))}_{(0)}$ and $W^{(one\_shot)}$ lie in the same sublevel set, while $W^{(min\_(0))}_{(0)}$ and $W^{(RPN\_2)}$ lie in different sublevel sets.
A comparison of top-100 positive eigen values of Hessians at $W^{(one\_shot)}$ and $W^{(RPN\_2)}$ is given in Fig. 18.
Figure 18: Comparison of top-100 positive eigen values of the Hessian at $W^{(one\_shot)}$ and $W^{(RPN\_2)}$.
It can be observed from the figure that $W^{(RPN\_2)}$ has larger eigen values of Hessian than that at $W^{(one\_shot)}$, which implies a smaller volume for the basin around $W^{(RPN\_2)}$ (Table VII).
TABLE VII: Comparison of $\lambda_{max}$ and $V^{{}^{\prime}}(100)$ at $W^{(RPN\_2)}$ and $W^{(one\_shot)}$.
Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
$W^{(RPN\_2)}$ 5635.883 487.644
$W^{(one\_shot)}$ 2944.881 450.409
@@
Fig. 19 presents a training loss along a straight line joining $W^{(min\_(9))}_{(9)}$ and $W^{(RPN\_1)}$. The figure shows a small barrier between $W^{(min\_(9))}_{(9)}$ and $W^{(RPN\_1)}$, indicating that they lie in the same sublevel set. However, the volume of the basin around $W^{(RPN\_1)}$ is smaller than the volume of the basin around $W^{(min\_(10))}_{(10)}$. This is illustrated in Fig. 20 and Table VIII.
Figure 19: Training Loss along a straight line between $W^{(min\_(9))}_{(9)}$ and $W^{(RPN\_1)}$.
Figure 20: Comparison of top-100 positive eigen values of the Hessian at $W^{(RPN\_1)}$ and $W^{(min\_(10))}_{(10)}$.
TABLE VIII: Comparison of $\lambda_{max}$ and $V^{{}^{\prime}}(100)$ at $W^{(RPN\_1)}$ and $W^{(min\_(10))}_{(10)}$.
Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
$W^{(RPN\_1)}$ 2100.56 435.624
$W^{(min\_(10))}_{(10)}$ 1575.91 408.81
@@ IV-G Result 7: Why fine-tuning doesn’t perform at par with rewinding?
Fig. 21 presents the comparison of top-100 positive eigen values of the Hessian at $W^{(FT)}$ and $W^{(min\_(10))}_{(10)}$.
Figure 21: Comparison of top-100 positive eigen values of the Hessian at $W^{(FT)}$ and $W^{(min\_(10))}_{(10)}$.
It can be observed from the figure that $W^{(FT)}$ has larger eigen values of the Hessian than $W^{(min\_(10))}_{(10)}$, which implies a smaller volume for the basin around $W^{(FT)}$ (Table IX). This elucidates the importance of rewinding the learning rate schedule while re-training the pruned network.
TABLE IX: Comparison of $\lambda_{max}$ and $V^{{}^{\prime}}(100)$ at $W^{(FT)}$ and $W^{(min\_(10))}_{(10)}$
. Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$ $W^{(FT)}$ 1667.231 412.686 $W^{(min\_(10))}_{(10)}$ 1575.910 408.810
@@ V Conclusion and Scope for Further Work
In this work, we have studied the loss landscape characteristics and volume/geometry of the IMP solutions at different levels in order to answer some important questions about the IMP and the lottery ticket hypothesis. The study provided the following important insights among others: 1) there exist a special type of minima in the loss landscape, which perform well but have a very small volume in the original space, and the IMP procedure uncovers such minima; 2) there exists a barrier between the IMP solutions at successive levels in the loss landscape; 3) IMP solutions obtained using rewinding lie within the same loss sublevel set, and random pruning or random initialization of a pruned network take the SGD out of that sublevel set. These insights help better understand the underlying principles behind IMP and the lottery-ticket hypothesis.
We have shown that there exist minima that have narrow profiles along certain dimensions and flatter profiles along others; however, the narrow profiles of these minima are along sparser dimensions. There is a possibility of the existence of minima that have narrow profiles along certain dimensions and flatter profiles along others, with narrow profiles along non-sparser dimensions. These minima are clearly not important from the pruning perspective. If so, an interesting question is: can such minima serve some other useful properties? This is a potential future direction that can be explored.
@@
Inverse volume of basin. In our experiments, we calculate the logarithm of the product of top-$100$ positive eigen values of the Hessian of the loss function to approximate the inverse volume of the basin. And the Hessian of the loss function is calculated on a randomly chosen subset of the training dataset ($1/5^{th}$ of the size of the full training dataset in the case of ResNet-20 and $1/10^{th}$ of the size of the full training dataset in the case of VGG-16).
@@
Loss landscape plot. Fig.10 presents the training loss at and around the neighbourhood of the different points of interest. The plot has been obtained by calculating training loss at $4200$ points in the high-dimensional space and then projecting the points on the two-dimensional plane spanned by $3$ points, $W_{(0)}^{(min\_(0))}$, $W_{(10)}^{(min\_(10))}$ and $W^{(RIPN)}$. However, the two axes formed by these points are not orthogonal. Hence, orthogonalization of these axes is carried out before projecting the points from the high-dimensional space. The loss is calculated using a randomly chosen subset of the training dataset ($1/5^{th}$ of the size of the full training dataset).
@@
Fig. 23 presents the comparison of training loss and test accuracy between $W_{(12)}^{(min\_(12))}$, $W^{(FT)}$, $W^{(RIPN)}$, $W^{(RPN)}$ and $W^{(one\_shot)}$. The plots clearly show that the solution obtained with IMP-WR ($W_{(12)}^{(min\_(12))}$) outperforms the solutions obtained with the other strategies like: fine-tuning ($W^{(FT)}$), random initialization of the pruned network ($W^{(RIPN)}$), random pruning ($W^{(RPN)}$) and one-shot pruning ($W^{(one\_shot)}$).
Figure 23: Comparison of training loss and test accuracy between $W_{(12)}^{(min\_(12))}$, $W^{(FT)}$, $W^{(RIPN)}$, $W^{(RPN)}$ and $W^{(one\_shot)}$ in case of VGG-16. Left: Training loss. Right: Test accuracy.
@@
Fig. 25 presents the comparison of top-100 positive eigen values of the Hessian at $W_{(L)}^{(min\_(L))}$ and $W^{Pr{(min\_(L-1))}}_{(L)}$ for $L=$$1,6,11$}. As before, the figure shows that the eigen values of the Hessian at $W_{(L)}^{(min\_(L))}$ are smaller than that at $W^{Pr{(min\_(L-1))}}_{(L)}$, which means larger volume for the basin around $W_{(L)}^{(min\_(L))}$ than the volume of the basin around $W^{Pr{(min\_(L-1))}}_{(L)}$. This is also depicted in Table X.
@@
TABLE X: Comparison of Maximum eigen value, $\lambda_{max}$ and inverse volume of basin, $V^{{}^{\prime}}(100)$ at $W^{Pr{(min\_(L-1))}}_{(L)}$ and $W_{(L)}^{(min\_(L))}$ for $L=$$1,6,11$}.
$L$ Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
1 $W^{Pr{(min\_(0))}}_{(1)}$ 70.099 212.779
$W_{(1)}^{(min\_(1))}$ 50.77 203.681
6 $W^{Pr{(min\_(5))}}_{(6)}$ 86.27 198.106
$W_{(6)}^{(min\_(6))}$ 37.08 177.233
11 $W^{Pr{(min\_(10))}}_{(11)}$ 1413.523 414.844
$W_{(11)}^{(min\_(11))}$ 1326.092 352.86
@@
Fig. 26 presents the comparison of top-100 positive eigen values of the Hessian at $W_{(L-1)}^{(min\_(L-1))}$ and $W^{RPr{(min\_(L))}}_{(L-1)}$ for $L=$$1,6,11$}. The figure shows that the eigen values of the Hessian at $W_{(L-1)}^{(min\_(L-1))}$ are smaller than that at $W^{RPr{(min\_(L))}}_{(L-1)}$, which means larger volume for the basin around $W_{(L-1)}^{(min\_(L-1))}$ than that of the basin around $W^{RPr{(min\_(L))}}_{(L-1)}$. This is also depicted in table XI.

### v3

@@ IV-C Result 3: Why do we need the Iterative Process, and why does one-shot pruning not work as well?
IV-D Result 4: There exists a barrier between IMP solutions at successive levels in the loss landscape.
@@
One interpretation of result 1 would clearly be that IMP solutions at different levels are distinct minima. In order to test this hypothesis more definitively, we examine whether these are separated by distinct barriers in the loss landscape. If this was not the case, SGD should converge to a sparser solution in the first place and would not stop at a less sparse solution. Evidence for our hypothesis is presented in Fig. 12. It gives the training loss along a straight line connecting $W^{(min\_(L-1))}_{(L-1)}$ and $W^{(min\_(L))}_{(L)}$. The barriers between successive minima are clearly visible. Thus, at least in the loss landscape, the IMP solutions at successive levels are not linearly connected, unlike the corresponding claim by [17] in the test error landscape. The implication is that these loss barriers are large enough to prevent SGD from crossing over and small enough to have almost similar generalization performance all over the region between these solutions.
@@
Figure 12: Training Loss along a straight line connecting $W^{(min\_(L-1))}_{(L-1)}$ and $W^{(min\_(L))}_{(L)}$ for $L$ ranging from $1$ to $10$. The x-axis represents the interpolation co-efficient $\alpha$. Each plot depicts the training loss at $501$ points between $W^{(min\_(L-1))}_{(L-1)}$ and $W^{(min\_(L))}_{(L)}$.
@@
IV-E Result 5: IMP solutions obtained using rewinding lie within the same loss sublevel set.
@@
Any random initialization of a dense neural network makes SGD converge to a good minimum. This can be ascribed to the well-understood fact that the density of good minima in such scenarios is very high [26]. However, in sparse subspaces, the density of such minima is much smaller. If we consider starting from a random point in such subspaces, the solution would not be as good as we get with rewinding, as is observed in our experiments.
@@
We also hypothesize that all the IMP solutions obtained using rewinding lie within the same loss sublevel set as that of the original dense network, whereas random initialization of the pruned network takes SGD out of that sublevel set. Fig. 13 presents the evidence for our claim. The plot has been obtained by calculating training loss at $4200$ points in the high-dimensional space and then projecting the points on the two-dimensional plane spanned by $3$ points, $W_{(0)}^{(min\_(0))}$, $W_{(10)}^{(min\_(10))}$ and $W^{(RIPN)}_{(10)}$. However, the two axes777 ($W_{(10)}^{(min\_(10))}-W_{(0)}^{(min\_(0))}$) forms the x-axis and ($W^{(RIPN)}_{(10)}-W_{(0)}^{(min\_(0))}$) forms the y-axis. formed by these points are not orthogonal. Hence, orthogonalization of these axes is carried out before projecting the points from the high-dimensional space. The figure clearly shows that all the IMP solutions ($W^{(min\_(0))}_{(0)}$ to $W^{(min\_(10))}_{(10)}$) lie within the same connected loss sublevel set (dark green region). However, $W^{(RIPN)}_{(10)}$ lies outside the sublevel set. Similarly, random pruning of a large number of weights also takes SGD outside the sublevel set. It is interesting to see that $W^{(one\_shot)}_{(10)}$ also lies in the same sublevel set as the IMP solutions; however, $W^{(RPN\_2)}_{(10)}$ lies outside the sublevel set. In other words, a randomly pruned network obtained by pruning a large number of parameters does not yield an equally good solution as that of the network obtained with magnitude based pruning, as it is not seen to lie in the same sublevel set.
@@
Figure 14: Euclidean distance between different points of interest. Left: Euclidean distance between IMP solutions at different levels. Right: Euclidean distance of $W^{(FT)}_{(10)}$, $W^{(RIPN)}_{(10)}$, $W^{(RPN\_1)}_{(10)}$, $W^{(one\_shot)}_{(10)}$ and $W^{(RPN\_2)}_{(10)}$ with $W^{(min\_(0))}_{(0)}$, $W^{(min\_(9))}_{(9)}$ and $W^{(min\_(10))}_{(10)}$
Figure 15: Cosine similarity between different points of interest. Left: Cosine similarity between IMP solutions at different levels. Right: Cosine similarity of $W^{(FT)}_{(10)}$, $W^{(RIPN)}_{(10)}$, $W^{(RPN\_1)}_{(10)}$, $W^{(one\_shot)}_{(10)}$ and $W^{(RPN\_2)}_{(10)}$ with $W^{(min\_(0))}_{(0)}$, $W^{(min\_(9))}_{(9)}$ and $W^{(min\_(10))}_{(10)}$.
It is apparent from these figures that the euclidean distance between IMP solutions is smaller than the euclidean distance of $W^{(RIPN)}_{(10)}$ with $W^{(min\_(9))}_{(9)}$ and $W^{(min\_(10))}_{(10)}$. Similarly, the cosine similarity between IMP solutions is greater than the cosine similarity of $W^{(RIPN)}_{(10)}$ with $W^{(min\_(9))}_{(9)}$ and $W^{(min\_(10))}_{(10)}$. Also the euclidean distance between $W^{(min\_(0))}_{(0)}$ and $W^{(one\_shot)}$ is smaller than the euclidean distance between $W^{(min\_(0))}_{(0)}$ and $W^{(RPN\_2)}_{(10)}$, while the cosine similarity between $W^{(min\_(0))}_{(0)}$ and $W^{(one\_shot)}_{(10)}$ is greater than the cosine similarity of $W^{(min\_(0))}_{(0)}$ with $W^{(RPN\_2)}_{(10)}$. This further buttresses the claim.
@@ IV-F Result 6: What happens when you prune the smaller weights?
$Loss({W_{pruned}})=Loss({W})+{(W_{pruned}-W)^{T}}\frac{\partial Loss}{\partial{W}}_{{W}={W}}$
$+{(W_{pruned}-W)^{T}}\frac{\partial^{2}Loss}{\partial{W}^{2}}{(W_{pruned}-W)}$ (6)
@@
Pruning out the smallest components in the weight vector ${W}$ results in only a small increase in the loss value. Pruning out the larger weights tends to bring about a bigger increase in the loss value since it makes the vector ${(W_{pruned}-W)}$ have a larger magnitude (norm). This can invalidate the second-order approximation given in Eq. $6$, since, in general, the loss function will not be a smooth one, and may have several points of inflection in the neighborhood.
@@
Randomly pruning a large number of weights also has a similar effect in that it may lead SGD to a position outside the sublevel set region, so that even re-training with weight rewinding cannot bring the network back to the sublevel set. Fig. 17 presents the training loss along a straight line connecting $W^{(min\_(0))}_{(0)}$ and $W^{(one\_shot)}_{(10)}$, and $W^{(min\_(0))}_{(0)}$ (baseline for $W^{(RPN\_2)}_{(10)}$) and $W^{(RPN\_2)}_{(10)}$.
Figure 17: Left: Training Loss along a straight connecting $W^{(min\_(0))}_{(0)}$ and $W^{(one\_shot)}_{(10)}$. Right: Training Loss along a straight connecting $W^{(min\_(0))}_{(0)}$ and $W^{(RPN\_2)}_{(10)}$.
The figure shows a small barrier between $W^{(min\_(0))}_{(0)}$ and $W^{(one\_shot)}_{(10)}$, and a huge barrier between $W^{(min\_(0))}_{(0)}$ and $W^{(RPN\_2)}_{(10)}$, which again demonstrates that $W^{(min\_(0))}_{(0)}$ and $W^{(one\_shot)}_{(10)}$ lie in the same sublevel set, while $W^{(min\_(0))}_{(0)}$ and $W^{(RPN\_2)}_{(10)}$ lie in different sublevel sets.
A comparison of top-100 positive eigen values of Hessians at $W^{(one\_shot)}_{(10)}$ and $W^{(RPN\_2)}_{(10)}$ is given in Fig. 18.
Figure 18: Comparison of top-100 positive eigen values of the Hessian at $W^{(one\_shot)}_{(10)}$ and $W^{(RPN\_2)}_{(10)}$.
It can be observed from the figure that $W^{(RPN\_2)}_{(10)}$ has larger eigen values of Hessian than that at $W^{(one\_shot)}_{(10)}$, which implies a smaller volume for the basin around $W^{(RPN\_2)}_{(10)}$.
@@
Fig. 19 presents a training loss along a straight line joining $W^{(min\_(9))}_{(9)}$ and $W^{(RPN\_1)}_{(10)}$. The figure shows a small barrier between $W^{(min\_(9))}_{(9)}$ and $W^{(RPN\_1)}_{(10)}$, indicating that they lie in the same sublevel set. However, the volume of the basin around $W^{(RPN\_1)}_{(10)}$ is smaller than the volume of the basin around $W^{(min\_(10))}_{(10)}$. This is illustrated in Fig. 20.
Figure 19: Training Loss along a straight line between $W^{(min\_(9))}_{(9)}$ and $W^{(RPN\_1)}_{(10)}$.
Figure 20: Comparison of top-100 positive eigen values of the Hessian at $W^{(RPN\_1)}_{(10)}$ and $W^{(min\_(10))}_{(10)}$.
@@ IV-G Result 7: Why fine-tuning doesn’t perform at par with rewinding?
Fig. 21 presents the comparison of top-100 positive eigen values of the Hessian at $W^{(FT)}_{(10)}$ and $W^{(min\_(10))}_{(10)}$.
Figure 21: Comparison of top-100 positive eigen values of the Hessian at $W^{(FT)}_{(10)}$ and $W^{(min\_(10))}_{(10)}$.
It can be observed from the figure that $W^{(FT)}_{(10)}$ has larger eigen values of the Hessian than $W^{(min\_(10))}_{(10)}$, which implies a smaller volume for the basin around $W^{(FT)}_{(10)}$. This elucidates the importance of rewinding the learning rate schedule while re-training the pruned network.
@@ V Conclusion and Scope for Further Work
In this work, we have studied the loss landscape characteristics and volume/geometry of the IMP solutions at different levels in order to answer some important questions about the IMP and the lottery ticket hypothesis. The study provided the following important insights among others: 1) there exist special type of solutions in the loss landscape, which perform well but have a very small volume in the original space, and the IMP procedure uncovers such solutions; 2) there exists a barrier between the IMP solutions at successive levels in the loss landscape; 3) IMP solutions obtained using rewinding lie within the same loss sublevel set, and random pruning or random initialization of a pruned network take the SGD out of that sublevel set. These insights help better understand the underlying principles behind IMP and the lottery-ticket hypothesis.
We have shown that solutions exist that have narrow profiles along certain dimensions and flatter profiles along others; however, the narrow profiles of these solutions are along sparser dimensions. There is a possibility of the existence of solutions that have narrow profiles along certain dimensions and flatter profiles along others, with narrow profiles along non-sparser dimensions. These solutions are clearly not important from the pruning perspective. If so, an interesting question is: can such solutions serve some other useful properties? This is a potential future direction that can be explored.
@@
Inverse volume of basin. In our experiments, we calculate the logarithm of the product of top-$100$ positive eigen values of the Hessian of the loss function to approximate the inverse volume of the basin. And the Hessian of the loss function is calculated on a randomly chosen subset of the training dataset ($1/5^{th}$ of the size of the full training dataset in the case of ResNet-20 and $1/10^{th}$ of the size of the full training dataset in the case of VGG-16). Fig. 22 shows that the trend with top-$200$ is consistent with top-$100$.
Figure 22: Comparison of top-200 positive eigen values of the Hessian at $W_{(10)}^{(min\_(10))}$ and $W^{Pr{(min\_(9))}}_{(10)}$ in case of ResNet-20.
@@
Loss landscape plot. Fig. 13 presents the training loss at and around the neighbourhood of the different points of interest. The loss is calculated using a randomly chosen subset of the training dataset ($1/5^{th}$ of the size of the full training dataset). The plot has been obtained by calculating training loss at $4200$ points in the high-dimensional space and then projecting the points on the two-dimensional plane spanned by $3$ points, $W_{(0)}^{(min\_(0))}$, $W_{(10)}^{(min\_(10))}$ and $W^{(RIPN)}$. However, the two axes formed by these points are not orthogonal. Hence, orthogonalization of these axes is carried out before projecting the points from the high-dimensional space.
Projection of a high-dimensional vector $r$ onto the plane spanned by orthogonal vectors $dx$ and $dy$ has been obtained as follows: $x=(r.dx)/|dx|$, $y=(r.dy)/|dy|$, where $x$ and $y$ are the projection coordinates along $dx$ and $dy$ respectively.
@@
In the case of VGG-16, one-shot pruned network is obtained by pruning the weights of the trained dense network $W_{(0)}^{(min\_(0))}$ based on magnitude pruning in one go to attain the desired sparsity (same sparsity as that of $W_{(12)}^{(min\_(12))}$) and then rewinding the unpruned weights to their values at $W^{(rewind\_point)}$ and retraining. We represent the solution obtained using one-shot pruning by $W^{(one\_shot)}_{(12)}$. The fine-tuned network is obtained by pruning $40\%$ smallest magnitude weights from $W_{(11)}^{(min\_(11))}$ and then re-training the unpruned weights (without rewinding) with a learning rate of $0.001$ for $40$ epochs. We represent the solution obtained using fine-tuning by $W^{(FT)}_{(12)}$. Randomly initialized pruned network is obtained by pruning $40\%$ smallest magnitude weights from $W_{(11)}^{(min\_(11))}$ and then randomly initializing the unpruned weights and retraining. We represent the solution by $W^{(RIPN)}_{(12)}$. A randomly pruned network has been obtained by randomly pruning $40\%$ weights from $W_{(11)}^{(min\_(11))}$ and then rewinding and retraining. We represent the solution obtained in the aforementioned manner by $W^{(RPN)}_{(12)}$. Fig. 24 presents the comparison of training loss and test accuracy between $W_{(12)}^{(min\_(12))}$, $W^{(FT)}_{(12)}$, $W^{(RIPN)}_{(12)}$, $W^{(RPN)}_{(12)}$ and $W^{(one\_shot)}_{(12)}$. The plots clearly show that the solution obtained with IMP-WR ($W_{(12)}^{(min\_(12))}$) outperforms the solutions obtained with the other strategies like: fine-tuning ($W^{(FT)}_{(12)}$), random initialization of the pruned network ($W^{(RIPN)}_{(12)}$), random pruning ($W^{(RPN)}_{(12)}$) and one-shot pruning ($W^{(one\_shot)}_{(12)}$).
Figure 24: Comparison of training loss and test accuracy between $W_{(12)}^{(min\_(12))}$, $W^{(FT)}_{(12)}$, $W^{(RIPN)}_{(12)}$, $W^{(RPN)}_{(12)}$ and $W^{(one\_shot)}_{(12)}$ in case of VGG-16. Left: Training loss. Right: Test accuracy.
@@
Fig. 26 presents the comparison of top-100 positive eigen values of the Hessian at $W_{(L)}^{(min\_(L))}$ and $W^{Pr{(min\_(L-1))}}_{(L)}$ for $L=$$1,6,11$}. As before, the figure shows that the eigen values of the Hessian at $W_{(L)}^{(min\_(L))}$ are smaller than that at $W^{Pr{(min\_(L-1))}}_{(L)}$, which means larger volume for the basin around $W_{(L)}^{(min\_(L))}$ than the volume of the basin around $W^{Pr{(min\_(L-1))}}_{(L)}$. This is also depicted in Table VI.
@@
TABLE VI: Comparison of inverse volume of basin, $V^{{}^{\prime}}(100)$ at $W^{Pr{(min\_(L-1))}}_{(L)}$ and $W_{(L)}^{(min\_(L))}$ for $L=$$1,6,11$}.
$L$ Solution $V^{{}^{\prime}}(100)$
1 $W^{Pr{(min\_(0))}}_{(1)}$ 212.779
$W_{(1)}^{(min\_(1))}$ 203.681
6 $W^{Pr{(min\_(5))}}_{(6)}$ 198.106
$W_{(6)}^{(min\_(6))}$ 177.233
11 $W^{Pr{(min\_(10))}}_{(11)}$ 414.844
$W_{(11)}^{(min\_(11))}$ 352.86
@@
Fig. 27 presents the comparison of top-100 positive eigen values of the Hessian at $W_{(L-1)}^{(min\_(L-1))}$ and $W^{RPr{(min\_(L))}}_{(L-1)}$ for $L=$$1,6,11$}. The figure shows that the eigen values of the Hessian at $W_{(L-1)}^{(min\_(L-1))}$ are smaller than that at $W^{RPr{(min\_(L))}}_{(L-1)}$, which means larger volume for the basin around $W_{(L-1)}^{(min\_(L-1))}$ than that of the basin around $W^{RPr{(min\_(L))}}_{(L-1)}$. This is also depicted in table VII.